OpenStudy (anonymous):

Find x if 5log2=logx

OpenStudy (lgbasallote):

\[5\log 2 = \log x\] so that means \[\implies \log 2^5 = \log x\] hint: if both sides have "log" then you can just cancel the logs out.....what do you get then?

OpenStudy (anonymous):

32?